# of the a an and or to from in as about but than then so through

### v1

@@ 5.4. Threats to validity
A first threat to validity concerns the generalizability of our findings. By sampling data from GitHub and carefully filtering it, we analyze real-world code. However, industrial and closed-source software may exhibit characteristics different from those observed in open-source projects on GitHub. Large parts of our methodology and implementation are reusable on other data sources or within industrial case studies, which could help assess the extent of these differences.
@@
Duplicate removal represents another potential threat. Our approach eliminates token-level duplicates from the dataset, but does not remove near-duplicates, i.e., files that differ only by a few characters or lines. Such near-duplicates can arise from different versions of the same library. Prior work by Lopes et al. (Lopes et al., 2017) addresses this issue using token-similarity thresholds. However, it remains unclear how to define an appropriate threshold or how to assess its validity in our setting. We leave the investigation of near-duplicate detection for future work.
@@
6. Discussion and Conclusion
@@
Although function sizes are comparable, our study shows that the internal structure and feature usage of FPBench functions differ from user-written code. In particular, transcendental functions are overrepresented, whereas conditionals and their combinations with other control-flow constructs are significantly underrepresented. FPBench thus aligns more closely with the capabilities of existing static floating-point reasoning techniques, which provide limited support (if any) for conditionals and loops  [^5] Handling conditionals in static analyses of e.g., rounding errors is challenging, because the real-valued specification and floating-point execution can take different paths through a program. Since worst-case rounding error bounds grow with every loop iteration, it is unclear how to find and prove loop invariants., and most techniques are not modular. Our study further reveals that, despite being frequently used for benchmarking in the literature, calls to GSL functions appear rarely (less than 0.2 %) at both the file and function level.
We believe that our corpus can be used to extract code for different benchmark sets. Future research can also rely on this study to better gauge whether the complexity in the benchmark accurately reflects real-world floating-point usage.
To be effective on user-written code, floating-point analysis tools must prioritize modularity and reasoning about branching behavior.
Our study shows that floating-point code is typically modular, with most function bodies containing multiple calls to others. Conditionals are common, and loops almost always co-occur with conditionals. These characteristics suggest that the main challenge for analysis tools is not handling large, monolithic numerical kernels, but effectively analyzing many small functions that combine modularity with non-trivial control flow. While tools such as PRECISA 4.0 (Titolo et al., 2025) and Hugo (Abbasi and Darulova, 2023) demonstrate promising approaches in this direction, the majority of state-of-the-art roundoff error analyzers target self-contained code fragments.
@@
This study presents part of the big picture, and extending the analysis to dynamically typed languages would provide a natural complement. However, we believe this cannot be achieved by simply adapting our current methodology. Our findings show that in statically typed code, non-type keywords appear only rarely (6 %-17 % of functions depending on the language), and GSL-related keywords appear even less frequently, making non-standard math libraries an unreliable proxy. Without explicit type annotations, a keyword-based approach in dynamically typed languages would therefore raise serious concerns about the validity of the results. The presence of arithmetic operators would also be insufficient to identify floating-point usage, since they are commonly overloaded for integers or non-numerical data. Obtaining reliable type information in such languages would require instrumenting and executing the code with representative inputs, which in turn demands resolving dependencies and reproducing realistic program executions (including GUI interactions where relevant). Designing methods to gather such information consistently and at scale remains an open challenge left for future work.

### v2

@@ 5.4. Threats to Validity
A first threat to validity concerns the generalisability of our findings. By sampling data from GitHub and carefully filtering it, we analyse real-world code. However, industrial, closed-source software or open-source software from other sources than GitHub may exhibit characteristics different from those observed in open-source projects on GitHub. Large parts of our methodology and implementation are reusable on other data sources or within industrial case studies, which could help assess the extent of these differences in future work.
@@
Duplicate removal represents another potential threat. Our approach eliminates token-level duplicates from the dataset, but does not remove near-duplicates, i.e., files that differ only by a few characters or lines. Such near-duplicates can arise from different versions of the same library. Prior work by Lopes et al. (lopesDejaVuMapCode2017) addresses this issue using token-similarity thresholds. However, it remains unclear how to define an appropriate threshold or how to assess its validity in our setting. We leave the investigation of near-duplicate detection for future work.
@@
6. Case Study: A Challenge Set of C Floating-Point Benchmarks
To demonstrate how our dataset can support practical benchmarking, we construct a set of 59 challenge benchmarks in C, which we upload along with the dataset. The choice to extract benchmarks in C is arbitrary (from the list of most frequently used languages for floating-point code). This case study is not intended as a definitive benchmark suite, but rather as an illustration of how our corpus can be used to generate realistic, ready-to-run benchmarks that reflect real-world floating-point practices. Each benchmark is a self-contained file containing a function extracted from our corpus together with all necessary dependencies to compile and run it independently.
Sampling and Extraction
The extraction process is semi-automated. We uniformly sample 200 functions from projects using permissive licenses (MIT, BSD, Apache-2.0). For each sampled function, we parse the source with Clang’s Rust bindings [^8] The extraction process is available as a subcommand in our framework. and recursively constructs its dependency graph by traversing the project’s directory structure. Dependencies include type definitions and auxiliary functions. When a new file is discovered, we also store the include directives and macros it contains. When all dependencies are identified, we emit a single self-contained C file that aggregates the collected includes and macros, followed by the function and its dependencies in topological order. Because we do not rely on build systems such as make, extraction can fail in the presence of conditional compilation or cyclic dependencies (46 out of 200 sampled functions). Manual curation removes redundant dependencies and macros. To preserve self-containment and portability, we also exclude benchmarks that depend on external or platform-specific libraries (e.g., OpenMP, CPython, windows.h), discarding 60 out of 155 remaining files. We further discard duplicates [^9] Although files in our dataset are deduplicated, identical function definitions may appear in distinct files, often due to code generation. (34 out of 94) and empty functions (1 out of 94), resulting in a final set of 59 unique benchmarks. We validate each extracted benchmark by compiling with both GCC 13.3.0 on Linux and Clang 21.1.8 on macOS.
double calSum(double data[], int N, int I) {
if (I == 0) { return data[0]; }
double sum = 0.f;
int idx = 0;
if (I >= N) { idx = N; }
else { idx = I; }
for (int i = 0; i < idx; i++) {
sum += data[I - i];
return sum;
void calAR(double PRICE[][4], int count, int N, double AR[]) {
double hoValue[count];
double olValue[count];
for (int i = 0; i < count; i++) {
hoValue[i] = PRICE[i][0] - PRICE[i][2];
olValue[i] = PRICE[i][2] - PRICE[i][1];
for (int i = 0; i < count; i++) {
double sum1 = calSum(hoValue, N, i);
double sum2 = calSum(olValue, N, i);
AR[i] = sum1 / sum2 * 100;
Figure 6. Example C floating-point benchmark extracted from our corpus. Unlike FPBench procedures, calAR is not self-contained and calls calSum (CalAR).
Two C functions, calSum and calAR, where calAR calls calSum to compute the sum of the first I elements of an array.
Characteristics
The resulting programs span multiple domains, including scientific computing (44 %), embedded systems (12 %), software toolchains (12 %), graphics (10 %), and educational material (10 %). Their size ranges from 4 to 1021 lines of code (LoC), with a median of 30 LoC (against 3 for FPBench). Unlike FPBench isolated numerical routines, the extracted functions preserve their original program context: several depend on auxiliary functions (29 %) or user-defined structures and unions (37 %), and most mix floating-point parameters with other types (66 %).
Figure˜6 illustrates one such benchmark, where the target function calAR calls an auxiliary routine calSum that computes the sum of the first I elements of an array. Benchmarks of this kind provide realistic cases for evaluating the interprocedural reasoning capabilities of floating-point analysis tools.
The extracted programs also differ from FPBench in the language constructs they employ. A large fraction use macros (31 %), pointer manipulation (63 %), and explicit type casts (41 %). They further exhibit floating-point usage patterns that extend beyond pure arithmetic or mathematical functions. For example, Figure˜7 shows a benchmark performing low-level memory manipulation to encode the least significant bits of an input into a NaN payload, while other programs use floating-point variables as loop counters. These patterns illustrate how, due to C’s permissive semantics, floating-point computations can appear in unexpected contexts. By exposing such patterns, these benchmarks prompt tool designers to make explicit and informed choices about which language features to support and which to explicitly exclude by design.
static float quiet_nanf(float x) {
uint32_t tmp;
memcpy(&tmp, &x, 4);
tmp |= 0x7fc00000lu;
memcpy(&x, &tmp, 4);
return x;
Figure 7. C floating-point benchmark extracted from our corpus that uses low-level memory manipulation to store the least significant bits of the input into a NaN value (quietNaN).
C function creating a quiet NaN by copying float bits to an integer, setting the NaN payload, and copying back.
By preserving real-world coding idioms that existing floating-point tools struggle to support, benchmarks extracted from our corpus will inevitably be challenging. We envision such benchmarks to serve as a challenging and realistic target for future tool development.
Table 7. Overview of representative freely available and recent floating-point reasoning tools. All listed tools support floating-point arithmetic.
[t] Tool input lang. aim supported features Static tools FPTaylor (Solovyev et al., 2018) custom worst-case rounding error bounds transcendental fncs. PRECiSA (Titolo et al., 2025) PVS worst-case rounding error bounds, mixed-precision tuning transcendental fncs., map and fold over lists, non-recursive func. callsa conditional stmt, loopsa Daisy (Abbasi and Darulova, 2023; Isychev and Darulova, 2023) Scala worst-case rounding error bounds transcendental fncs., some vector and matrix operations, non-recursive func. calls Gappa (Daumas and Melquiond, 2010) custom worst-case rounding error bounds 11 rounding modes (only arithm.) Frama-C (Boldo and Marché, 2011) C deductive program verification essentially full C, limited support for rounding errors via Gappa KeY (Abbasi et al., 2023) Java deductive program verification subset of Java (roughly basic Java 1.2), no autom. support for rounding errors Stainless (Gilot et al., 2026) Scala deductive program verification subset of Scala (mostly functional), no autom. support for rounding errors Dynamic tools Herbie (Panchekha et al., 2015) S-expressions improve accuracy via rewriting conditionals, fixed set of mathematical functions Verrou (Févotte and Lathuilière, 2017) integrated to Valgrind numerical instability via Monte Carlo arithmetic presumably all Verificarlo (Denis et al., 2016) C, C++, Fortran numerical instability via Monte Carlo arithmetic presumably all FPLearner (Wang and Rubio-González, 2024b) C/C++ mixed-precision tuning presumably all NSan (Courbet, 2021) LLVM large rounding errors via shadow execution presumably all a Not evaluated systematically in publications as far as we know.
Applicability to Existing Tools
Table˜7 provides an overview of a selection of existing tools. This list is not and is not meant to be exhaustive, rather we focus on available and recent tools that could potentially benefit from a benchmark suite such as the one that we extracted. (As stated previously, we do not intend our extracted C benchmarks as a definitive benchmark suite.) We focus on recent tools, i.e. tools whose code was updated in some way within the last 5 years, as older tools tend to be difficult to run in practice. The threshold of 5 years is arbitrary.
Many dynamic analysis tools target C/C++ in some way and so we expect that they can be immediately run on the extracted benchmarks. To use the tool Herbie, one would need to extract the numerical calculations out of the benchmarks.
To demonstrate that our benchmarks can be used in practice, we run NSan and Frama-C with Gappa as frontend on them. In contrast to other tools, NSan requires an explicit entry point with a main function. To accommodate this requirement, we add a simple main function to each benchmark that invokes the extracted function with fixed input values chosen to avoid runtime errors. Frama-C does not require an entry point and we therefore run it directly on the extracted benchmarks. NSan successfully analyses all the benchmarks and does not report precision-loss warnings (this is not very surprising, as such warnings are typically highly input dependent). Frama-C produces verification conditions for 54 out of the 59 benchmarks; the remaining five benchmarks use complex numbers, which are not supported by Frama-C.
For the remainder of the static analysis tools, one would need to translate them from C to their respective input language. We expect that as part of this rewrite, some of the currently unsupported features of each tool (e.g., pointers) could also be expressed in a way that the tool can handle. This effort is beyond the scope of this paper.
An alternative option for these tools is to extract programs from GitHub directly written in the input language of interest. The newly added support for floating-point arithmetic in Stainless (Gilot et al., 2026) successfully used Scala programs extracted in this code study for evaluation.
7. Discussion and Conclusions
@@
Although function sizes are comparable, our study shows that the internal structure and feature usage of FPBench functions differ from user-written code. In particular, transcendental functions are overrepresented, whereas conditionals and their combinations with other control-flow constructs are significantly underrepresented. FPBench thus aligns more closely with the capabilities of existing static floating-point reasoning techniques, which provide limited support (if any) for conditionals and loops  [^10] Handling conditionals in static analyses of e.g., rounding errors is challenging, because the real-valued specification and floating-point execution can take different paths through a program. Since worst-case rounding error bounds grow with every loop iteration, it is unclear how to find and prove loop invariants., and whose techniques are not modular. Our study further reveals that, despite being frequently used for benchmarking in the literature, calls to GSL functions appear rarely (less than 0.2 %) at both the file and function level.
We believe that our corpus can be used to extract code for different benchmark sets. Future research can also rely on this study to better gauge whether the complexity in a benchmark accurately reflects real-world floating-point usage.
To be effective on user-written code, floating-point analysis tools must prioritize modularity and reasoning about branching behaviour.
Our study shows that floating-point code is typically modular, with most function bodies containing multiple calls to others. Conditionals are common, and loops almost always co-occur with conditionals. These characteristics suggest that the main challenge for analysis tools is not handling large, monolithic numerical kernels (Das et al., 2020), but effectively analysing many small functions that combine modularity with non-trivial control flow. While tools such as PRECISA 4.0 (Titolo et al., 2025) and Daisy (Abbasi and Darulova, 2023) demonstrate promising approaches in this direction, the majority of state-of-the-art roundoff error analysers target self-contained code fragments.
@@
This study presents part of the big picture, and extending the analysis to dynamically typed languages would provide a natural complement. However, we believe this cannot be achieved by simply adapting our current methodology. Our findings show that in statically typed code, non-type keywords appear only rarely (6 %–17 % of functions depending on the language), and GSL-related keywords appear even less frequently, making non-standard math libraries an unreliable proxy. Without explicit type annotations, a keyword-based approach in dynamically typed languages would therefore raise serious concerns about the validity of the results. The presence of arithmetic operators would also be insufficient to identify floating-point usage, since they are commonly overloaded for integers or non-numerical data. Obtaining reliable type information in such languages would require instrumenting and executing the code with representative inputs, which in turn demands resolving dependencies and reproducing realistic program executions (including GUI interactions where relevant). Designing methods to gather such information consistently and at scale remains an open challenge left for future work.
Semantically characterising floating-point code in a quantitative, scalable, and automated fashion raises substantial challenges.
While it would be very interesting to identify a repository’s application domain(s), detect common numerical algorithms, or assess numerical stability, these characterisations typically require manual annotation with multiple annotators to yield meaningful quantitative claims—we are not aware of automated tools that can do this reliably. Type-based analyses are similarly not as easily automated since they generally require successfully building projects. Prior work has attempted this for Java using build tools such as Maven or Gradle (sulirQuantitativeStudyJava2016; hassanAutomaticBuildingJava2017; sulirLargeScaleDatasetLocal2020), but reported low success rates, which can bias the analysis. The process described in Section˜6 to compile C benchmarks required several weeks of work for a single programming language and required manual intervention for a large number of projects. Developing a pipeline that addresses these issues is left for future work and can build on the baseline established by this paper.
8. Data-Availability Statement
All code, datasets, and artifacts necessary to reproduce the analyses and results presented in this paper are publicly available and permanently archived.
  • Scyros. The source code of Scyros, our framework for large-scale code studies, is available on GitHub under the Apache 2.0 license at: https://github.com/fxpl/scyros. A versioned snapshot corresponding to the evaluated artifact is archived on Zenodo along with the artifact.
  • Function Dataset. The corpus of metadata for floating-point functions extracted from GitHub, including links to the original functions, is archived on Zenodo under the Apache License 2.0 at: https://doi.org/10.5281/zenodo.17055622.
  • C Challenge Benchmarks: the set of 59 self-contained C floating-point benchmarks extracted from our corpus is available on Zenodo as part of the function dataset.
  • Floating-Point Keywords List: the list of keywords we used to identify floating-point code is also available on Zenodo as part of the function dataset.
  • Artifact. The artifact accompanying this paper is a local web application that reproduces all analyses and results reported in the paper and provides a step-by-step guide for reproducing the study with different keywords on a small set of newly mined repositories. The artifact is archived on Zenodo under the Apache 2.0 license at: https://doi.org/10.5281/zenodo.18500268.
Acknowledgments
We thank the reviewers for their thoughtful feedback that strengthened our work. Additionally, we would like to thank Fridtjof Stoldt, Noé De Santo and Paulo Canelas for their suggestions and feedback on this work.
This work is co-funded by the European Union (ERC, HORNET, 101163629). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
This research was supported by the Sponsor Swedish Research Council https://vr.se through the grant Grant #2023-05161, “Scalable Error Verification: How Accurate is your Numerical Code?”, and Grant #2024-04565, “Data-race Freedom and Memory Safety for Untyped Languages”.